OpenStudy (anonymous):

Use an infinite series to evaluate the definite integral of x^2e^(-x^2) from 0 to 1. The above problem looks like something I've done before, but I'm having trouble finding it in my notes. Perhaps someone could give me a clue as to how to get started?

OpenStudy (anonymous):

\[\int\limits_{0}^{1} x^2e^{-x^2}\] For those having trouble understanding the question as I typed it. I'm supposed to use an infinite series to evaluate it.

ganeshie8 (ganeshie8):

Recall the infinite series definition of e^x : \[\large e^x = \sum \limits_{n=0}^{\infty }\dfrac{x^n}{n!}\]

ganeshie8 (ganeshie8):

plugin x = -x^2, multiply both sides by x^2, then take the integral

OpenStudy (anonymous):

Why do I multiply both sides by x^2?

ganeshie8 (ganeshie8):

look at the integral you're trying to evaluate

ganeshie8 (ganeshie8):

the idea is to cookup that integral from this series representation of e^x

OpenStudy (anonymous):

Ah. Okay. So when I integrate both sides, do I still include sigma in the notation, or am I only taking the integral of x^2(-x^2)^n/n! on the right side?

ganeshie8 (ganeshie8):

when the series converges : \[\large \int \sum = \sum \int \]

ganeshie8 (ganeshie8):

\[\large e^x = \sum \limits_{n=0}^{\infty }\dfrac{x^n}{n!}\] \[\large e^{-x^2} = \sum \limits_{n=0}^{\infty }\dfrac{(-x^2)^n}{n!} = \sum \limits_{n=0}^{\infty }\dfrac{(-1)^nx^{2n}}{n!} \] \[\large x^2e^{-x^2} = \sum \limits_{n=0}^{\infty }\dfrac{(-1)^nx^{2n+2}}{n!} \]

ganeshie8 (ganeshie8):

integrate both sides

ganeshie8 (ganeshie8):

\[\large \int_0^1x^2e^{-x^2}dx = \int_0^1\sum \limits_{n=0}^{\infty }\dfrac{(-1)^nx^{2n+2}}{n!} dx\]

ganeshie8 (ganeshie8):

\[\large = \sum \limits_{n=0}^{\infty } \int_0^1 \dfrac{(-1)^nx^{2n+2}}{n!} dx \]

ganeshie8 (ganeshie8):

\[\large = \sum \limits_{n=0}^{\infty } \dfrac{(-1)^n}{n!} \int_0^1 x^{2n+2} dx \]

ganeshie8 (ganeshie8):

whats the integral of \(\large \int_0^1 x^{2n+2} dx \) ?

OpenStudy (anonymous):

\[\int\limits_{0}^{1} x^{2n+2} = \frac{ x^{2n+3} }{ 2n+3 } evaluated from 0 \to 1\]

ganeshie8 (ganeshie8):

yes evaluate the bounds

OpenStudy (anonymous):

After evaluating the bounds, you get \[\frac{ (1)^{2n+3} }{ 2n+3 }\] n doesn't have a value with regards to the integral, so I assume you multiply that result with the infinite series to get: \[\sum_{0}^{\infty} \frac{ (-1)^n(1)^{2n+3} }{ n!(2n+3) }\]

ganeshie8 (ganeshie8):

that looks good ^^ but hey (1)^anything = 1 right ?

OpenStudy (anonymous):

Right! So, after that it would simplify to: \[\sum_{0}^{\infty} \frac{ (-1)^n }{ n!(2n+3) }\]

ganeshie8 (ganeshie8):

Perfect !

OpenStudy (anonymous):

And I'm guessing all I need now to finish the evaluation is to find the sum of the infinite series and I'm done?

ganeshie8 (ganeshie8):

You're done! leave it as infinite series...

OpenStudy (anonymous):

Cool! Thanks.

ganeshie8 (ganeshie8):

yw!